OpenStudy (anonymous):

Your friend is having difficulty applying the chain rule for differentiation. Using the example g(t)=cos^2(x^3+x^2) , explain how to apply the chain rule to differentiate g(t)

OpenStudy (kagıtucak):

first of all you need to reduce the order then differantiate the inside the cos so 2.cos(x^3+x^2).(3x^2+2x).(-)sin(x^3+x^2)

zepdrix (zepdrix):

Is the function supposed to be g(x) ?

OpenStudy (anonymous):

probably

OpenStudy (kagıtucak):

oh i didnt realize that if it is g(t) solution will change

OpenStudy (anonymous):

if it doesnt change the solution, im sure whether its x or t doesnt matter. if it does change, keep it at g(t)

zepdrix (zepdrix):

Chain rule can be kind of tricky. I would try to get the "friend" to understand that a function consists of `layers`. `functions within functions`. Our outermost function is \(\Large g(x)=x^2\) If we stuff the function \(\Large h(x)=\cos x\) into g(x) we get,\[\Large g(h(x))=\left[\cos(x)\right]^2=\cos^2x\]If we then stuff the function \(\Large j(x)=x^3+x^2\) inside of g(h(x)) we get,\[\Large g(h(j(x)))=\cos^2(x^3+x^2)\] Then we just apply the chain rule as the rule tells us:\[\Large \frac{d}{dx}g(h(j(x)))\quad=\quad g'(h(j(x)))\cdot h'(j(x))\cdot j'(x)\] Hmm I think I just made it more confusing :( But if you want to be technical, and explain it using the definition for chain rule, then this might help...

OpenStudy (anonymous):

oh and its not actually my friend its an open ended promt

zepdrix (zepdrix):

that's why I said "friend" hehe

OpenStudy (anonymous):

ok and im not exactly sure what im supposed to do. its open ended so i assume i have to explain it using the definition and mathematically. its kind of vague.

zepdrix (zepdrix):

Let,\[\Large g(x)=x^2\]\[\Large h(x)=\cos x\]\[\Large j(x)=x^3+x^2\]Then,\[\Large f(x)\quad=\quad g(h(j(x)))\]and by the definition of chain rule,\[\Large f'(x)\quad=\quad g'(h(j(x)))\cdot h'(j(x))\cdot j'(x)\]And then list out these derivatives individually,\[\Large g'(h(j(x)))\quad=\quad 2\left[\cos(x^3+x^2)\right]\]\[\Large h'(j(x))\quad=\quad -\sin(x^3+x^2)\]\[\Large j'(x)\quad=\quad 3x^2+2x\]So our final answer f'(x) is the product of these three thingies. So we just need to multiply them together. Mmmmmmm I dunno D: That's how I would approach it maybe. Blah I kinda did all the work for you.. sorry bout that... takes the fun away.

OpenStudy (anonymous):

i multiplied them together and got: -2(3x^2+2x)sin(x^3+x^2)cos(x^3+x^2). would that be g'(t)?

zepdrix (zepdrix):

Ahhh crap, i was calling the original function f(x). I didn't notice that they labeled it as g(x). Sorry about that.. That makes things a little confusing lol. But yes, that would be your .... f'(x) or g'(x) as they labeled it. But not g'(t). I think that was a typo, hopefully..

OpenStudy (anonymous):

ok thanks!

zepdrix (zepdrix):

If it IS supposed to be g'(t), then we would have one more `chain`. We would throw a \(\Large \dfrac{dx}{dt}\) on the end.

OpenStudy (anonymous):

ok